# A noise-robust Monte Carlo method for electric field calculations in EMC3

### v1

@@ Abstract
One of the main codes to analyze and optimize stellarator configurations is the EMC3 code, which implements a state-of-the-art 3D Monte Carlo plasma edge transport code. However, so far, a self-consistent treatment of the $\bm{E}\times\bm{B}$ drift is absent. This plasma drift is known to significantly impact the particle and heat distribution in the plasma edge. It is desirable to incorporate this drift into EMC3 to improve the predictive capabilities of the code. The calculation of the $\bm{E}\times\bm{B}$ drift requires the approximation of the electric field $\bm{E}$, which is proportional to the gradient of the electric potential $\varphi$. In previous work [2], the gradient was calculated with a least squares method based on a finite difference approximation of the electric potential. However, due to the stochastic nature of EMC3, the output plasma fields computed by the code are inherently noisy. The finite difference method further amplifies the noise, with the amplification growing as the grid size decreases. We continue from [1], which introduced a new noise-robust method for 1D derivatives. We extend the noise-robust method to 2D and apply it to the electric potential. We show that a PDE can be derived that describes the evolution of the electric field in case of a uniform diffusion coefficient. This PDE allows us to approximate the electric field directly with a Monte Carlo simulation, thus avoiding the need for a finite difference approximation. We illustrate the accuracy of the method and the noise robustness with a test case.
@@ 1 Introduction
Nuclear fusion reactors have been a research topic for decades now. Over the years, countless of machine design were developed [6]. The most popular machine desgins are the tokamak and the stellerator. The stellerator omits the need of an induced plasma current to generate the poloidal magnetic field [7]. Instead, it relies on geometrically complex magnetic coils to induce spiraling magnetic field lines. Therefore, the mathematical models and corresponding numerical solvers have to deal with this complex 3D geometry.
A popular code to perform simulations of the plasma edge in stellerator type devices is the EMC3-EIRENE code. The MC-MC coupled code estimates the density, momentum and energy of the plasma. The plasma edge simulations require an efficient and accurate computation of gradients. Important gradients include the temperature gradient, pressure gradient, gradient of the electrical potential and the magnetic field gradient [3, 9, 10]. Espescially the gradient of the electical potential $\varphi$ plays an important role in plasma edge simulations, as it directly determines the electric field $\bm{E}$, through the relation $\bm{E}=-\nabla\varphi$. The electric field, in turn, is essential for modeling the $\bm{E}\times\bm{B}$ drift, which is a key mechanism for improving the predictive capabilities of the EMC3 code [4]. The main objective is therefore to obtain an accurate representation of the electric field $\bm{E}$ by reliably approximating the gradient of the electric potential $\varphi$. The electrical potential itself is approximated by the EMC3 code.
@@
In Section 2, we present the electric potential equation. The current approach to calculate the electric field uses a MC method to solve the electrical potential equation and then applies the FD method. Next, Section 3 introduces the Monte Carlo Gradient Approximation method where we derive a new equation that governs the evolution of the electric field directly. We additionaly highlight the challenges associated with this equation. Section 4 provides a numerical experiment to illustrate the ability and accuracy of the MCGA method. We also analyse the effect of grid resolution on the variance of both MCGA and FD approximations. Finally, Section 5 summarizes the main findings and presents the conclusions.
@@ 2 Electric Potential Equation
In this section, we derive a reduced model for the electrostatic potential $\varphi$. Charge conservation requires the current density $\mathbf{J}$ to be divergence-free, i.e.
@@
where $\sigma_{||}$ is the parallel conductivity, $p_{e}$ and $T_{e}$ denote the electron pressure and temperature, and $n_{e}$ is the electron density.
@@
$\mathbf{J_{\perp}}=\mathbf{J}_{\text{drift}}-\sigma_{\perp}^{an}\nabla_{\perp}\varphi,$ (4)
with $\sigma_{\perp}^{an}$ representing the anomalous perpendicular conductivity.
Substituting these expressions into Eq. (2) yields a diffusion equation for the electrostatic potential,
@@
$S=-\nabla\cdot\bm{J}_{\text{drift}}-\nabla\cdot\left(\frac{\sigma_{\perp}}{e}\left(T_{e}\nabla_{\parallel}\ln(n)+1.71\nabla_{\parallel}T_{e}\right)\right).$ (6)
@@
are employed. Other boundary condition, such as reflective boundaries, are also possible. Equation (5) is a stationary diffusion-type equation. However, in Sections 3 and 4, we consider a more general time-depenent diffusion equation. For clarity and simplicity, we adopt classical Dirichlet boundary conditions; however, the approach remains valid if Bohm boundary conditions are used instead. Equation (5) serves as the starting point for deriving a closed evolution equation for $\bm{E}$ in the next section.
@@ 3 Monte Carlo Gradient Approximation
In this section, we introduce the Monte Carlo Gradient Approximation method for computing gradients in 2D. For both the derivation and the numerical experiments, the electrical potential $\varphi$ is modeled by a time-dependent linear diffusion equation with spatially varying isotropic diffusion coefficient $D(\bm{r})$ in a Cartesian coordinate system. The equation equals
@@
where $\bm{r}=\left(x,y\right)^{T}$ and $\nabla=\left(\partial_{x},\partial_{y}\right)^{T}$. The quantity of interest is the electric field $\bm{E}=-\nabla\varphi$. The idea of the Monte Carlo Gradient Approximation method is to derive an evolution equation directly for $\bm{E}$. Taking the gradient of Equation (8) and assuming sufficient smoothness of $\varphi$ and $D$ to interchange the order of differentiation, we obtain
@@
Next, we swap the order of $\nabla$ and $\partial_{t}$, and susbstitute $\bm{E}=-\nabla\varphi$. This results in
@@
We aim to solve Equation (11) with a Monte Carlo method. However, at present time the writers are not familiar with any Monte Carlo method that can directly approximate vector quantities. Instead, we decompose Eq. (11) into scalar equations for the components $E_{x}$ and $E_{y}$
@@
We illustrate how these equations are coupled. Consider the source term of the $x$-component
@@
The underlined term couples different components of the electric field through the source term, which poses difficulties. The term $\partial_{y}E_{y}$ is unknown, and cannot be incorporated in the simulation in a straightforward way. Estimating $\partial_{y}E_{y}$ with finite differences would reintroduce the original problem we seek to resolve. In this work, the term $\partial_{y}E_{y}$ is neglected in numerical experiments, assuming that it introduces a small bias in the simulation. Further work is required to develop a consistent approximation strategy including these source terms.
@@ 4.1 Numerical Verification
As a first test, we validate the Monte Carlo method on a simple 2D problem on a square domain. We consider the following equation for the electrical potential:
@@
This differs from the electrical potential equation in Section 2 because Eq. (14) is time-dependent, the diffusion coefficient is isotropic, and the source term (6) is neglected. To improve readability, we omit the explicit function arguments in the following equations. We employ the method of manufactured solutions to determine the source term $S(t,x,y)$. We propose an exact solution $\varphi(x,y,t)=\exp(-(x+y))\exp(-Dt)$, from which the source term follows as
@@
The corresponding exact electrical field components are
$E_{x}(t,x,y)=-\exp(-(x+y))\exp(-Dt),$ $E_{y}(t,x,y)=-\exp(-(x+y))\exp(-Dt).$ (16)
Differentiating Equation (14) with respect to $x$ and $y$ yields the governing equations for the electric field components. For $E_{x}$, we obtain
@@
$S_{E_{x}}=\left(2D-4Dx\right)\varphi+\left(-D+2Dx-2Dx^{2}\right)E_{x}+\underline{2Dx\partial_{y}E_{y}}.$ (18)
@@
Dirichlet boundary conditions are imposed for $\varphi,E_{x}$ and $E_{y}$ with boundary values calculated from the exact solution. The initial condition is likewise taken from the exact solution evaluated in $t=t_{0}$. All three equations (14),(17) and (19) are simulated with identical parameter values, namely
@@
where $N$ denotes the number of particles per simulation, $t_{0}$ the simulation start time, $T^{*}$ the end time and $\Delta t$ the time step. The domain is $\left[0,1\right]\times\left[0,1\right]$ discretized with $M_{x}=M_{y}=15$ grid points. In the source term (18), we utilize the Monte Carlo approximation of $\varphi$. The MC approximation of $\varphi$ is computed jointly with $E_{x}$ and $E_{y}$, such that three equations are solved in parallel. The source terms are incorporated using an operator-splitting method [5]. A detailed description on the treatment of source terms can be found in [1].
Regarding the underlined term in Equation (18), two approaches are considered. Since the manufactured solution is known, $\partial_{y}E_{y}$ can be evaluated exactly. When an exact solution is missing, this term is set to 0 under the assumption that it has a minor contribution to the dynamics.
Figure 1 presents the electrical field $\bm{E}$ and the norm $\|\bm{E}\|_{2}$, approximated using both the MCGA method and a first-order FD method. For the MCGA method, we additionaly show the results obtained when $\partial_{y}E_{y}$ is either set to zero or given by the exact source term. The Monte Carlo approximation exhibits significantly less noise compared to the finite difference approximation. Additonally, neglecting the term $\partial_{y}E_{y}$ does not introduce a significant bias in the approximation, likely because in this test case the term contributes negligibly to the overall dynamics.
@@
Figure 2 shows the relative error of the MCGA method, both when neglecting $\partial_{y}E_{y}$ and when the exact term is used, and the relative error of the FD method. The FD method displays a consistently higher relative error across the domain. When $\partial_{y}E_{y}$ is neglected, the relative error is of the same order of magnitude and the approximation remains equally accurate.
@@ 4.2 Noise Robustness
Next, we investigate the effect of grid resolution on the variance of the MCGA method, and compare the results with the FD method. While a Monte Carlo method is independent of the grid size, the grid resolution affects the density approximation. In 2D, for a uniform grid and equal scaling in both dimensions, the variance increases with order 2. For the FD method, we expect an increase of order 4, consistent with Eq. (1).
To avoid complications arising from incomputable source terms, we consider the simplified diffusion problem for the electrical potential
$\partial_{t}\varphi(t,x,y)=D\nabla^{2}\left(\varphi(t,x,y)\right),$ (22)
@@
$E_{x}(t,x,y)=-\frac{2x}{4\pi Dt}\exp(-\frac{x^{2}+y^{2}}{4Dt}),$ $E_{y}(t,x,y)=-\frac{2y}{4\pi Dt}\exp(-\frac{x^{2}+y^{2}}{4Dt}),$ (24)
@@
$\partial_{t}E_{x}(t,x,y)=D\nabla^{2}(E_{x}(t,x,y)),$ $\partial_{t}E_{y}(t,x,y)=D\nabla^{2}(E_{y}(t,x,y)),$ (25)
respectively. Equations (22)-(25) are solved on the domain $\left[0,2\right]\times\left[0,2\right]$ using uniform grids with a different number of gridpoints $M=M_{x}=M_{y}=\{11,21,41,81\}$. The remaining simulation parameters are chosen as
@@
where $\hat{u}_{M}$ denotes the solution, calculated with either MC or FD, in the middle of the grid with $M$ grid points. We ran 20 independent simulations to approximate the expectation values with Welford’s online algorithm [11]. The results are displayed in Figure 3. We observe the theoretical predicted order of convergence. Furthermore, they indicate that the MC method exhibits greater robustness with respect to grid refinement than the FD method.
Figure 3: Effect of the grid resolution on the variance. The variance of the FD approximation of $E_{x}$ grows faster than the MC approximation of $E_{x}$. This indicates that the MC method is more robust to noise increase with decreasing grid sizes.
@@ 5 Conclusion
In this work, we presented the Monte Carlo Gradient Approximation method for computing the electrical field. We demonstrated how to derive a new equation for the electical field from the electrical potential and discussed the challenges associated with approximating the source term in this formulation. With a numerical experiment, we illustrated that the method works and achieves a lower relative error than FD, even when we set the $\partial_{y}E_{y}$ to zero. Lastly, we showed that the method achieves better robustness to noise compared to finite differences for decreasing grid sizes.
For future work, more realistic test cases need to be considered to further assess the accuracy and applicability of the Monte Carlo Gradient Approximation method. These may include cases with Bohm boundary conditions and non-rectangular grids. Such test cases are necessary to assess whether neglecting $\partial_{y}E_{y}$ is justified. If omitting this source term introduces a significant bias, more advanced strategies for handling source terms in the electric field equations are needed.
@@
William De Deyn: Conceptualizations (lead), Writing software, Writing of the original draft. Ruben De Wolf: Conceptualizations (support), Helping in writing of Section 2, Review and editing. Vince Maes: Conceptualizations (support), Review and editing. Giovanni Samaey: Conceptualizations (support), Review and editing.

### v2

@@ Abstract
One of the main codes to analyze and optimize stellarator configurations is the EMC3 code, which implements a state-of-the-art 3D Monte Carlo plasma edge transport code. However, so far, a self-consistent treatment of the $\bm{E}\times\bm{B}$ drift is absent. This plasma drift is known to significantly impact the particle and heat distribution in the plasma edge. It is desirable to incorporate this drift into EMC3 to improve the predictive capabilities of the code. The calculation of the $\bm{E}\times\bm{B}$ drift requires the approximation of the electric field $\bm{E}$, which is proportional to the gradient of the electric potential $\varphi$. In previous work [2], the gradient was calculated with a least squares method based on a finite difference approximation of the electric potential. However, due to the stochastic nature of EMC3, the output plasma fields computed by the code are inherently noisy. The finite difference method further amplifies the noise, with the amplification growing as the grid size decreases. We build on [1], which introduced a new noise-robust method for 1D derivatives. We extend the noise-robust method to 2D and apply it to the electric potential. We show that a PDE can be derived that describes the evolution of the electric field in the case of an isotropic diffusion coefficient. This PDE allows us to approximate the electric field directly with a Monte Carlo simulation, thus avoiding the need for a finite difference approximation. We illustrate the accuracy of the method and the noise robustness with a test case.
@@ 1 Introduction
Nuclear fusion reactors have been a research topic for decades now. Over the years, countless machine design were developed [6]. The most popular machine designs are the tokamak and the stellarator. The stellarator eliminates the need for an induced plasma current to generate the poloidal magnetic field [7]. Instead, it relies on geometrically complex magnetic coils to induce spiraling magnetic field lines. Therefore, the mathematical models and corresponding numerical solvers have to deal with this complex 3D geometry.
A popular code to perform simulations of the plasma edge in stellarator type devices is the EMC3-EIRENE code. The MC-MC coupled code estimates the density, momentum, and energy of the plasma. The plasma edge simulations require an efficient and accurate computation of gradients. Important gradients include the temperature gradient, pressure gradient, gradient of the electric potential, and the magnetic field gradient [3, 9, 10]. Especially the gradient of the electric potential $\varphi$ plays an important role in plasma edge simulations, as it directly determines the electric field $\bm{E}$, through the relation $\bm{E}=-\nabla\varphi$. The electric field, in turn, is essential for modeling the $\bm{E}\times\bm{B}$ drift, which is a key mechanism for improving the predictive capabilities of the EMC3 code [4]. The main objective, therefore, is to obtain an accurate representation of the electric field $\bm{E}$ by reliably approximating the gradient of the electric potential $\varphi$. The electric potential itself is approximated by the EMC3 code.
@@
In Section 2, we present the electric potential equation. The current approach to calculate the electric field uses a MC method to solve the electric potential equation and then applies the FD method. Next, Section 3 introduces the Monte Carlo Gradient Approximation method where we derive a new equation that governs the evolution of the electric field directly. We additionally highlight the challenges associated with this equation. Section 4 provides a numerical experiment to illustrate the ability and accuracy of the MC method. We also analyze the effect of grid resolution on the variance of both MC and FD approximations. Finally, Section 5 summarizes the main findings and presents the conclusions.
@@ 2 Electric Potential Equation
In this section, we derive a reduced model for the electric potential $\varphi$. Charge conservation requires the current density $\mathbf{J}$ to be divergence-free, i.e.
@@
where $\sigma_{\parallel}$ is the parallel conductivity, $p_{e}$ and $T_{e}$ denote the electron pressure and temperature, and $n_{e}$ is the electron density.
@@
$\mathbf{J_{\perp}}=\mathbf{J}_{\text{drift}}-\sigma_{\perp}\nabla_{\perp}\varphi,$ (4)
with $\sigma_{\perp}$ representing the anomalous perpendicular conductivity.
Substituting these expressions into Eq. (2) yields a diffusion equation for the electric potential,
@@
$S=-\nabla\cdot\bm{J}_{\text{drift}}-\nabla\cdot\left(\frac{\sigma_{\parallel}}{e}\left(T_{e}\nabla_{\parallel}\ln(n_{e})+1.71\nabla_{\parallel}T_{e}\right)\right).$ (6)
@@
are employed. Other boundary conditions, such as reflective boundaries, are also possible. Equation (5) is a stationary diffusion-type equation. However, in Sections 3 and 4, we consider a more general time-dependent diffusion equation. For clarity and simplicity, we adopt classical Dirichlet boundary conditions; however, the approach remains valid if Bohm boundary conditions are used instead. Equation (5) serves as the starting point for deriving a closed evolution equation for $\bm{E}$ in the next section.
@@ 3 Monte Carlo Gradient Approximation
In this section, we introduce the Monte Carlo Gradient Approximation method for computing gradients in 2D. For both the derivation and the numerical experiments, the electric potential $\varphi$ is modeled by a time-dependent linear diffusion equation with spatially varying isotropic diffusion coefficient $D(\bm{r})$ in a Cartesian coordinate system. The equation is
@@
where $\bm{r}=\left(x,y\right)^{\top}$ and $\nabla=\left(\partial_{x},\partial_{y}\right)^{\top}$. The quantity of interest is the electric field $\bm{E}=-\nabla\varphi$. The idea of the Monte Carlo Gradient Approximation method is to derive an evolution equation directly for $\bm{E}$. Taking the gradient of Equation (8) and assuming sufficient smoothness of $\varphi$ and $D$ to interchange the order of differentiation, we obtain
@@
Next, we swap the order of $\nabla$ and $\partial_{t}$, and substitute $\bm{E}=-\nabla\varphi$. This results in
@@
We aim to solve Equation (11) with a Monte Carlo method. However, at present, we are not aware of any Monte Carlo method that can directly approximate vector quantities. Instead, we decompose Eq. (11) into scalar equations for the components $E_{x}$ and $E_{y}$
@@
We illustrate how these equations are coupled. Consider the source term for the $x$-component
@@
The underlined term couples different components of the electric field through the source term, which poses difficulties. The term $\partial_{y}E_{y}$ is unknown, and cannot be incorporated into the simulation in a straightforward way. Estimating $\partial_{y}E_{y}$ with finite differences would reintroduce the original problem we seek to resolve. In this work, the term $\partial_{y}E_{y}$ is neglected in numerical experiments, assuming that it introduces a small bias in the simulation. Further work is required to develop a consistent approximation strategy to include these source terms.
@@ 4.1 Numerical Verification
As a first test, we validate the Monte Carlo method on a simple 2D problem on a square domain. We consider the following equation for the electric potential:
@@
This differs from the electric potential equation in Section 2 because Eq. (14) is time-dependent, the diffusion coefficient is isotropic, and the source term (6) is neglected. To improve readability, we omit the explicit function arguments in the following equations. We employ the method of manufactured solutions to determine the source term $S(t,x,y)$. We propose an exact solution $\varphi(x,y,t)=\exp(-(x+y))\exp(-Dt)$, from which the source term follows as
@@
The corresponding exact electric field components are
$E_{x}(t,x,y)=\exp(-(x+y))\exp(-Dt),$ $E_{y}(t,x,y)=\exp(-(x+y))\exp(-Dt).$ (16)
Differentiating Equation (14) with respect to $x$ and $y$, and substituting $E_{x}=\partial_{x}\varphi$ and $E_{y}=\partial_{y}\varphi$, yields the governing equations for the electric field components. For $E_{x}$, we obtain
@@
$S_{E_{x}}=\left(-2D+4Dx\right)\varphi+\left(-D+2Dx-2Dx^{2}\right)E_{x}+\underline{2Dx\partial_{y}E_{y}}.$ (18)
@@
Dirichlet boundary conditions are imposed for $\varphi,E_{x}$, and $E_{y}$ with boundary values calculated from the exact solution. The initial condition is likewise taken from the exact solution evaluated at $t=t_{0}$. All three equations (14), (17), and (19) are simulated with identical parameter values, namely
@@
where $N$ denotes the number of particles per simulation, $t_{0}$ the simulation start time, $T^{*}$ the end time and $\Delta t$ the time step. The domain is $\left[0,1\right]\times\left[0,1\right]$ discretized with $M_{x}=M_{y}=15$ grid points. In the source term (18), we utilize the Monte Carlo approximation of $\varphi$. The MC approximation of $\varphi$ is computed jointly with $E_{x}$ and $E_{y}$, such that three equations are solved in parallel. The source terms are incorporated using an operator-splitting method [5]. A detailed description of the treatment of source terms can be found in [1].
Regarding the underlined term in Equation (18), two approaches are considered. Since the manufactured solution is known, $\partial_{y}E_{y}$ can be evaluated exactly. When an exact solution is missing, this term is set to zero under the assumption that it has a minor contribution to the dynamics.
Figure 1 presents the electric field $\bm{E}$ and the norm $\|\bm{E}\|_{2}$, approximated using both the MCGA method and a first-order FD method. For the MCGA method, we additionally show the results obtained when $\partial_{y}E_{y}$ is either set to zero or given by the exact source term. The Monte Carlo approximation exhibits significantly less noise compared to the finite difference approximation. Additionally, neglecting the term $\partial_{y}E_{y}$ does not introduce a significant bias in the approximation, likely because in this test case the term contributes negligibly to the overall dynamics.
@@
Figure 2 shows the relative error of the MCGA method, both when neglecting $\partial_{y}E_{y}$ and when using the exact term, as well as the relative error of the FD method. The FD method displays a consistently higher relative error across the domain. When $\partial_{y}E_{y}$ is neglected, the relative error is of the same order of magnitude and the approximation remains equally accurate.
@@ 4.2 Noise Robustness
Next, we investigate the effect of grid resolution on the variance of the MCGA method, and compare the results with the FD method. While a Monte Carlo method is independent of the grid size, the grid resolution affects the density approximation. In 2D, for a uniform grid and equal scaling in both dimensions, the variance increases quadratically. For the FD method, we expect an increase of order 4, consistent with Eq. (1).
To avoid complications arising from intractable source terms, like $\partial_{y}E_{y}$, we consider the simplified diffusion problem for the electric potential
$\partial_{t}\varphi(t,x,y)=D\nabla^{2}\varphi(t,x,y),$ (22)
@@
$E_{x}(t,x,y)=\frac{2x}{4\pi Dt}\exp(-\frac{x^{2}+y^{2}}{4Dt}),$ $E_{y}(t,x,y)=\frac{2y}{4\pi Dt}\exp(-\frac{x^{2}+y^{2}}{4Dt}),$ (24)
@@
$\partial_{t}E_{x}(t,x,y)=D\nabla^{2}E_{x}(t,x,y),$ $\partial_{t}E_{y}(t,x,y)=D\nabla^{2}E_{y}(t,x,y),$ (25)
respectively. Equations (22)-(25) are solved on the domain $\left[0,2\right]\times\left[0,2\right]$ using uniform grids with different number of grid points $M=M_{x}=M_{y}=\{11,21,41,81\}$. The remaining simulation parameters are chosen as
@@
where $\hat{u}_{M}$ denotes the solution, calculated with either MC or FD, in the middle of the grid with $M$ grid points. We ran 20 independent simulations to approximate the expectation values with Welford’s online algorithm [11]. The results are displayed in Figure 3. We observe the theoretically predicted order of convergence. Furthermore, they indicate that the MC method exhibits greater robustness with respect to grid refinement than the FD method.
Figure 3: Effect of the grid resolution on the variance. The variance of the FD approximation of $E_{x}$ grows faster than the MC approximation of $E_{x}$. This indicates that the MC method is more robust to noise increases in noise as the grid size decreases.
@@ 5 Conclusion
In this work, we presented the Monte Carlo Gradient Approximation method for computing the electric field. We demonstrated how to derive a new equation for the electrical field from the electric potential and discussed the challenges associated with approximating the source term in this formulation. With a numerical experiment, we illustrated that the method works and achieves a lower relative error than FD, even when we set $\partial_{y}E_{y}$ to zero. Lastly, we showed that the method achieves better robustness to noise compared to finite differences for decreasing grid sizes.
For future work, more realistic test cases need to be considered to further assess the accuracy and applicability of the Monte Carlo Gradient Approximation method. These may include cases with Bohm boundary conditions and non-rectangular grids. Such test cases are necessary to assess whether neglecting $\partial_{y}E_{y}$ is justified. If omitting this source term introduces a significant bias, more advanced strategies for handling source terms in the electric field equations may be needed.
@@
William De Deyn: Conceptualization (lead), Software implementation, Writing of the original draft. Ruben De Wolf: Conceptualization (support), Writing Section 2 (support), Review and editing. Vince Maes: Conceptualization (support), Review and editing. Giovanni Samaey: Conceptualization (support), Review and editing.